The molecular masses of organo-oxygen or organosulfur (chalcogens) compounds were found to be:

a) 108 amu

b) 122 amu

i. What is a possible chemical formula for the compound for each case?

ii. What is the units of unsaturation in each case?
Answer #1

a)

i)

Given the molecular mass of a organo-oxygen molecule is 108 amu

The possible organo-oxygen compound with the given molecular mass is : C7H8O

Since the atomic mass of C is 12 amu, H is 1 amu and O is 16 amu.

So total molecular mass becomes : (7x12) amu + (1x8) amu + (1x16) amu = (84 amu + 8 amu + 16 amu )= 108 amu

The possible organosulfur compound with the given molecular mass of 108 amu is : C6H4S

Since the atomic mass of C is 12 amu, H is 1 amu and S is 32 amu.

So total molecular mass becomes : (6x12) amu + (4x1) amu + (1x32) amu = (72 amu + 4 amu + 32 amu )= 108 amu

ii) Degree of unsaturation of a compound with molecular formula CAHBXCODNE is :

(A+1) - (B/2) - (C/2) + (E/2) --------------------- (1)

Use this formula (1), to find the degree of unsaturation in C7H8O:

DU = (7+1) - (8/2)

= 8-4

= 4

hence, the units of unsaturation is 4.

Use this formula (1), to find the degree of unsaturation in C6H4S:

DU = (6+1) - (4/2)

= 7-2

= 5

b)

i)

Given the molecular mass of a organo-oxygen molecule is 122 amu

The possible organo-oxygen compound with the given molecular mass is : C8H10O

Since the atomic mass of C is 12 amu, H is 1 amu and O is 16 amu.

So total molecular mass becomes : (8x12) amu + (1x10) amu + (1x16) amu = (96 amu + 10 amu + 16 amu )= 122 amu

The possible organosulfur compound with the given molecular mass of 108 amu is : C7H6S

Since the atomic mass of C is 12 amu, H is 1 amu and S is 32 amu.

So total molecular mass becomes : (7x12) amu + (6x1) amu + (1x32) amu = (84 amu + 6 amu + 32 amu )= 122 amu

ii)

Degree of unsaturation of a compound with molecular formula CAHBXCODNE is :

(A+1) - (B/2) - (C/2) + (E/2) --------------------- (1)

Use this formula (1), to find the degree of unsaturation in C8H10O:

DU = (8+1) - (10/2)

= 9-5

= 4

Use this formula (1), to find the degree of unsaturation in C7H6S:

DU = (7+1) - (6/2)

= 8-3

= 5

Degree of unsaturation represents the number of double bonds or number of triple bonds or the number of rings present in the given compound.